# MetAdv: A Unified and Interactive Adversarial Testing Platform for Autonomous Driving

### v1

@@ 1. Introduction
Autonomous driving has made significant progress in recent years (Zimmer et al., 2019; Lang et al., 2019; Erçelik et al., 2022; Zimmer et al., 2022; Li et al., 2024; Hu et al., 2023b; Chen et al., 2024; Jiang et al., 2023), yet remain highly susceptible to adversarial attacks (Szegedy et al., 2013; Goodfellow et al., 2014), ranging from digital perturbations in sensor input (Dong et al., 2023; Hallyburton et al., 2022; Liu et al., 2020b; Zhang et al., 2021) to physical-world camouflage (Eykholt et al., 2018; Liu et al., 2019; Wang et al., 2021; Liu et al., 2021, 2020a, 2023a, 2022, 2023c; Guo et al., 2023; Liu et al., 2023b; Zhang et al., 2024b; Wang et al., 2024, 2025; Zhang et al., 2024a), which can lead to critical failures and safety hazards. As these systems move closer to widespread deployment, evaluating and ensuring their adversarial robustness has become an urgent and unresolved challenge. However, current adversarial testing approaches are hampered by the lack of a unified, interactive, and dynamic evaluation platform. Existing studies (Xu et al., 2022; Li et al., 2022; Tang et al., 2021) often focus narrowly on isolated and static environments, limiting their capacity to uncover the full spectrum of weaknesses that may emerge in real-world conditions.

### v2

@@ 1. Introduction
Autonomous driving has made significant progress in recent years (Zimmer et al., 2019; Lang et al., 2019; Erçelik et al., 2022; Zimmer et al., 2022; Li et al., 2024; Hu et al., 2023b; Chen et al., 2024; Jiang et al., 2023), yet remain highly susceptible to adversarial attacks (Szegedy et al., 2013; Goodfellow et al., 2014; Liang et al., 2020; Wei et al., 2018; Liang et al., 2022b, a; Kong et al., 2024), ranging from digital perturbations in sensor input (Dong et al., 2023; Hallyburton et al., 2022; Liu et al., 2020b; Zhang et al., 2021) to physical-world camouflage (Eykholt et al., 2018; Liu et al., 2019; Wang et al., 2021; Liu et al., 2021, 2020a, 2023a, 2022, 2023c; Guo et al., 2023; Liu et al., 2023b; Zhang et al., 2024b; Wang et al., 2024, 2025; Zhang et al., 2024a), which can lead to critical failures and safety hazards. As these systems move closer to widespread deployment, evaluating and ensuring their adversarial robustness has become an urgent and unresolved challenge. However, current adversarial testing approaches are hampered by the lack of a unified, interactive, and dynamic evaluation platform. Existing studies (Xu et al., 2022; Li et al., 2022; Tang et al., 2021) often focus narrowly on isolated and static environments, limiting their capacity to uncover the full spectrum of weaknesses that may emerge in real-world conditions.